OpenStudy (anonymous):

Factor: (3x-4)^2 + (x-5)(2)(3x-4)(3)

OpenStudy (mayliz):

(3x-4)^2 + (x-5)(2)(3x-4)(3)=x (27 x-138)+136 =27 x^2-138 x+136

OpenStudy (anonymous):

Oh my gosh!! That's what I got!!! :D I thought it was supposed to be left with stuff in parenthesis!!!

OpenStudy (anonymous):

Oh, I see! You can't break it down anymore!! Thank you so much!!!!!